OpenStudy (anonymous):

how to factor 6x^2+7x+2 ???

OpenStudy (cwrw238):

you need 2 numbers whose product is 6*2 (12) and sum is 7

OpenStudy (cwrw238):

can you think of them?

OpenStudy (cwrw238):

how about 3 and 4 you need to get 7x by 3x + 4x :- 6x^2+7x+2 = 6x^2 + 3x + 4x + 2 = 3x(2x + 1) + 2(2x + 1) = 3x + 2(2x + 1)

OpenStudy (cwrw238):

* (3x + 2)(2x + 1)

OpenStudy (goformit100):

"Welcome to OpenStudy. I can guide regarding this useful site; ask your doubts from me, for it you can message me. Please use the chat for off topic questions. And remember to give the helper a medal, by clicking on "Best Answer". We follow a code of conduct, ( http://openstudy.com/code-of-conduct ). Please take a moment to read it."